LONG-WAVELENGTH PHOTOCHEMISTRY OF MATRIX-ISOLATED BIACETYL

\underline{NATHAN G. KUCHMAS} and C. A. BAUMANN, Department of Chemistry, The University of Scranton, Scranton, PA 18510-4626.

Visible irradiation (520 nm> λ >485 nm) of matrix-isolated biacetyl (C_{4}H_{6}O_{2}) results in the formation of a complex of cis-methylhydroxycarbene (CH_{3}COH) and ketene (CH_{2}CO), as well as other products. The wavelengths used in this study are longer than those necessary for the S_{0}-S_{1} (X̃^{1}A_{g} -Ã^{1}A_{u}) transition, indicating that the photolysis is the result of a multiphoton process. One such process is sequential, where the spin-forbidden S_{0}-T_{1} (X̃^{1}A_{g} -ã^{3}A_{u}) transition is followed by the allowed T_{1}-T_{2} (ã^{3}A_{u} -b̃^{3}B_{g}) transition, with photoproducts emerging from the higher triplet (or following a radiationless transition to another state). A simultaneous two-photon process is also possible through the symmetryforbidden S_{0}-S_{2} (X̃^{1}A_{g} -B̃^{1}B_{g}) transition. Photoproduct formation as a function of irradiation flux and wavelength was used to sort out the relative contributions of each of these processes.